Circular No. 7943

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ef IN NGC 7761 B. Beutler and W. D. Li, University of California at Berkeley, report

the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on unfiltered KAIT CCD images taken on July 30.4 (mag ∼ 16.5) and 31.4 UT (mag ∼ 16.2). SN 2002ef is located at α = 23^{h}51^{m}29^{s}.54, δ = −13^{o}22′46′′.4 (equinox 2000.0), which is 10′′.0 east and 7′′.0 north of the nucleus of NGC 7761. A KAIT image of the same field on July 16.4 UT showed nothing at this position (limiting mag ∼ 18.0).

V838 MONOCEROTIS H. E. Bond, N. Panagia, and W. B. Sparks, Space Telescope Science

Institute; S. G. Starrfield, Arizona State University; R. M. Wagner, University of Arizona; and A. Henden, Universities Space Research Association and U.S. Naval Observatory, report that the light echo around V838 Mon was imaged with the Hubble Space Telescope and Advanced Camera for Surveys on May 20. An empty hole in the center of the light echo has grown slightly in size since the previous ACS observation on Apr. 30 (cf. IAUC 7892). Preliminary analysis of images taken in the V band plus polarizing filters reveals a narrow ring of highly (∼ 20%) linearly polarized light, centered on the star and lying just outside the inner edge of this hole at a radius of about 4′′.8. The high polarization indicates that the light scattering occurs at an angle of nearly 90^{o}, and thus the material producing it lies close to the distance of the star itself. This implies a purely geometric lower limit to the stellar distance of 2.5 kpc, if the polarized ring is illuminated by light emitted at the March 11 peak of the star’s V band outburst. The outermost rings in the light echo are at a radius of 20′′, corresponding to material lying about 0.5 pc in front of the star, and are expanding as the square root of the time since outburst. The central hole suggests that there is a cavity in the dust surrounding the star, with a radius of about 2 light-months. If this cavity is spherical, we predict that light from the outburst will soon reach the back side of the cavity, thus converting the hole into a bright, filled center. Continued observations are urged when V838 Mon emerges from conjunction with the Sun.

SUPERNOVA 2002ed IN NGC 5468 H. Yamaoka, Kyushu University, Japan, reports the following posi-

tion for SN 2002ed from the discovery images (cf. IAUC 7940): α = 14^{h}06^{m}38^{s}.22, δ = −5^{o}27′28′′.0 (equinox 2000.0).

2002 July 31 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT Carl W. Hergenrother